Please give solutions with detailed explanations. I appreciate neat hand writing. Thanks a lot.3. (a) Show that [e-21/a (1 + 2r/a + 2r2/a?)] = -4r²e-20/a. (b) Use this result to show that the ground state wavefunction of
Answer #1

9a*(1+84 +132)] - * (2 +13) **(1+2+ 3 *** (379-A--) -- nem 6) Hydrogen ground state care function . 1 as Bohr rative t = e lawe have her an zu* (1+24+*%) = 0, hay altopical rule. Štai pag ds. 07(2* (****) Hence ground state of hydrogen atom is proper

menores The Bohir vadive is 0.529 Ao (ii) we have energy of hylogen atom is En = - 13-6 eV. for is, n-i, F, = -136 eV for as,but frequency 2=% hc = DE h = 4.136x16 els Ca 3x105 nls - DE -158 hc = 4.136X10 X 3x10 DE 10.2 = 1.2165x10m - 121.65 nm in th